# Lithium Niobate Photonic Topological Insulator-based Multi-Wavelength Optical Demultiplexer with Piezoelectric Switch-Off

### v1

@@ 4 Conclusion
In conclusion, a valley photonic topologically insulating 2-channel demultiplexer based on a lithium niobate photonic crystal with a wide range of operating wavelengths and a piezoelectric switch-off mechanism has been designed and demonstrated in this work. The existence of edge states, robust topological transport along the interface, optical insulation at the bulk, and immunity to backscattering, sharp turns, and defects was verified. A simple linear relation was derived between the operating wavelength and the structure’s lattice constant. Topological protection was preserved for each case of wavelength shift, thus introducing a simple method of designing the topological structure for multiple wavelength regions. An optical demultiplexer was formed using our design, which was shown to separate = 1310 nm and = 1550 nm components in isolated channels with good bandwidth. Topological insulation allows the crosstalk to be as low as 54 dB. Lastly, a method of piezoelectric switching of the topological protection was demonstrated, which can decrease the optical transmittance to less than 5 with a 2 V external potential. Other possible applications of our design include a piezoelectric optical filter, etc. This topological design will be helpful in applications that require lossless, robust optical transport, separation of wavelength components, and electrical switching of the optical transmittance. It will contribute to lithium niobate photonics, efficient optical communication systems, and photonic integrated circuits.

### v3

@@ 4 Conclusion
In conclusion, a valley photonic topologically insulating 2-channel demultiplexer based on a lithium niobate photonic crystal with a wide range of operating wavelengths and a piezoelectric switch-off mechanism has been designed and demonstrated in this work. The existence of edge states, robust topological transport along the interface, optical insulation at the bulk, and immunity to backscattering, sharp turns, and defects was verified. A simple linear relation between the operating wavelength and the lattice constant of the structure was derived. Topological protection was preserved for each case of wavelength shift, thus introducing a simple method of designing the topological structure for multiple wavelength regions. An optical demultiplexer was formed using our design, which was shown to separate = 1310 nm and = 1550 nm components in isolated channels with good bandwidth. Topological insulation allows the crosstalk to be as low as 54 dB. Lastly, a method of piezoelectric switching of the topological protection was demonstrated, which can decrease the optical transmittance to less than 5 with a 2 V external potential. Other possible applications of our design include a piezoelectric optical filter, etc. This topological design will be helpful in applications that require lossless, robust optical transport, separation of wavelength components, and electrical switching of the optical transmittance. It will contribute to lithium niobate photonics, efficient optical communication systems, and photonic integrated circuits.